# Human Mesh Recovery from Arbitrary Multi-view Images

### v1

@@ Abstract
Human mesh recovery from arbitrary multi-view images involves two characteristics: the arbitrary camera poses and arbitrary number of camera views. Because of the variability, designing a unified framework to tackle this task is challenging. The challenges can be summarized as the dilemma of being able to simultaneously estimate arbitrary camera poses and recover human mesh from arbitrary multi-view images while maintaining flexibility. To solve this dilemma, we propose a divide and conquer framework for Unified Human Mesh Recovery (U-HMR) from arbitrary multi-view images. In particular, U-HMR consists of a decoupled structure and two main components: camera and body decoupling (CBD), camera pose estimation (CPE), and arbitrary view fusion (AVF). As camera poses and human body mesh are independent of each other, CBD splits the estimation of them into two sub-tasks for two individual sub-networks (i.e., CPE and AVF) to handle respectively, thus the two sub-tasks are disentangled. In CPE, since each camera pose is unrelated to the others, we adopt a shared MLP to process all views in a parallel way. In AVF, in order to fuse multi-view information and make the fusion operation independent of the number of views, we introduce a transformer decoder with a SMPL parameters query token to extract cross-view features for mesh recovery. To demonstrate the efficacy and flexibility of the proposed framework and effect of each component, we conduct extensive experiments on three public datasets: Human3.6M, MPI-INF-3DHP, and TotalCapture.
@@ 3.1 Human Body Representation
The SMPL (Skinned Multi-Person Linear Model) model [37] is a skinned vertex-based model and can accurately represent a wide variety of body shapes in natural human poses. It provides a function $M(\theta,\beta;\Phi):\mathbb{R}^{|\theta|\times|\beta|}\mapsto\mathbb{R}^{3N}$, mapping pose and shape parameters to $N=6,890$ vertices to form a mesh representation of human body. In the function $\theta\in\mathbb{R}^{24\times 3}$, $\beta\in\mathbb{R}^{10}$, and $\Phi$ are pose, shape, and model parameters, respectively. The pose parameters $\theta$ includes global rotation $\theta_{g}\in\mathbb{R}^{3}$ which is unrelated to the body pose and body pose parameters $\theta_{b}\in\mathbb{R}^{23\times 3}$, consequently $\theta=\{\theta_{g},\theta_{b}\}$. The shape parameters $\beta$ is coefficient of a low-dimensional shape space, learned from a training set of thousands of registered scans.
Typically, a pre-trained linear regressor $W$ is used to obtain $k$ keypoints of interest of the whole body, i.e., $J_{3D}\in\mathbb{R}^{k\times 3}=WM(\theta,\beta;\Phi)$ [24]. As for 2D keypoints, a perspective camera model with fixed focal length and intrinsics $K$ is used to project 3D joint positions into 2D image plane. Each camera pose $\pi=\{R,t\}$ consists of a global orientation $R\in\mathbb{R}^{3}$ and a translation $t\in\mathbb{R}^{3}$. Given above parameters, 3D keypoints can be projected to 2D image as $J_{2D}=\Pi(K(RJ_{3D}+t))$, where $\Pi$ is a perspective projection with camera intrinsics $K$. Since the pose parameters $\theta$ already includes a global orientation $\theta_{g}$ which can represent the camera rotation, in practice we consider $R=\theta_{g}$, so when reconstructing human body from a single image, we estimate body pose parameters $\theta_{b}$ and camera pose $\pi=\{R,t\}$, and when reconstructing human body from $N$ images, we estimate body pose parameters $\theta_{b}$ and $N$ camera poses ${\pi_{i}=\{R_{i},t_{i}\}}_{i=1}^{N}$.
@@ 3.2 Unified Human Mesh Recovery (U-HMR)
Essentially, the multi-view feature fusion is a process of weighted sum of each view feature. In the transformer decoder structure, because of the cross-attention mechanism, the number of the weights is not hard coded in network, it is varying with the number of tokens. Due to this online variability, we can fuse features from arbitrary multi-view images. In addition, benefiting from the feature representation learning ability of the transformer decoder, the the updated SMPL query token can be more representative for human mesh recovery.
@@ 4.1 Experimental Setup
Human3.6M is a large 3D human pose benchmark, containing 11 different subjects and each subject with 15 different daily indoor actions. The subjects are captured by 4 synchronized 50 Hz digital cameras. The dataset provides 2D/3D keypoint annotations obtained by marker-based motion capture system. The SMPL annotations are obtained by applying MoSh [36] to the sparse 3D MoCap marker following previous works [20, 25]. Similar to [20, 31, 47], we follow the protocol 1 to use subjects 1, 5, 6, 7, and 8 for training and subjects 9 and 11 are for testing.
MPI-INF-3DHP contains 8 actors (4 males and 4 females) performing 8 human activities. 2D and 3D keypoint annotations are also provided by marker-less motion capture system. The dataset contains multi-view images of 14 cameras with regular lenses, and we choose views 0, 2, 7, 8 from them. Following [31, 47], the standard test dataset of MPI-INF-3DHP simply includes single-view images, we use subjects 1-7 for training and subject 8 for testing.
@@ 4.2 Ablation Studies
Table 2: Ablation study of number of views on Human3.6M dataset. Note that PaFF [19] uses camera calibrations for fusion, and the results reported here are from the models re-trained using 2 and 3 views respectively, while Yu et al. [61] and our method require neither camera calibrations nor retraining, so the results of PaFF are expected to provide the performance upper bound. We use ResNet-50 as the 2D image encoder for our method as the other two methods do for the fair of comparison.
@@
As we claim that our method can reconstruct human pose and shape for arbitrary multi-view images, to validate the flexibility, i.e., the ability to adapt to different number of views, we conduct ablation study on the effect of varying number of views during inference phase. Since there have been two methods, Yu et al. [61] and PaFF [19], state that they can be applied to arbitrary number of views, we compare with them on Human 3.6M dataset and the results are shown in Sec. 4.2. For fair comparison, our model is trained with a combination of varying number of multi-view image (1-view, 2-view, 3-view, and 4-view), which is same as  [61] does. While PaFF claims to have the ability to be applied to different number of views, the reported results are from the model retrained and tested using same number of views. Moreover, PaFF requires camera calibrations for feature fusion. So the results of PaFF can be expected to provide the performance upper bound of other two methods which are calibration-free and not retrained. Compared with Yu et al. [61], our method achieves better performance at any number of views.
@@ 4.3 Comparison with State-of-the-art Methods
Table 3: Comparison results on Human3.6M dataset. We compare our method with single-view, calibration-requiring multi-view, and calibration-free multi-view methods. ${}^{\dagger}$ denotes the output of methods are non-parametric.
@@
4.4 Computational Overhead
@@ 4.5 Sensitivity to Input
Table 7: Sensitivity to input. We perturb the centers and scales of ground truth bboxs in Human3.6M dataset and test model without retraining.
@@
Table 7: Sensitivity to input. We perturb the centers and scales of ground truth bboxs in Human3.6M dataset and test model without retraining.
5 Conclusion

### v4

@@ Abstract
Human mesh recovery from arbitrary multi-view images involves two characteristics: the arbitrary camera poses and arbitrary number of camera views. Because of the variability, designing a unified framework to tackle this task is challenging. The challenges can be summarized as the dilemma of being able to simultaneously estimate arbitrary camera poses and recover human mesh from arbitrary multi-view images while maintaining flexibility. To solve this dilemma, we propose a divide and conquer framework for Unified Human Mesh Recovery (U-HMR) from arbitrary multi-view images. In particular, U-HMR consists of a decoupled structure and two main components: camera and body decoupling (CBD), camera pose estimation (CPE), and arbitrary view fusion (AVF). As camera poses and human body mesh are independent of each other, CBD splits the estimation of them into two sub-tasks for two individual sub-networks (i.e., CPE and AVF) to handle respectively, thus the two sub-tasks are disentangled. In CPE, since each camera pose is unrelated to the others, we adopt a shared MLP to process all views in a parallel way. In AVF, in order to fuse multi-view information and make the fusion operation independent of the number of views, we introduce a transformer decoder with a SMPL parameters query token to extract cross-view features for mesh recovery. To demonstrate the efficacy and flexibility of the proposed framework and effect of each component, we conduct extensive experiments on three public datasets: Human3.6M, MPI-INF-3DHP, and TotalCapture. Code is available at: https://github.com/XiaobenLi00/U-HMR.
@@ 3.1 Human Body Representation
The SMPL (Skinned Multi-Person Linear Model) model [37] is a skinned vertex-based model and can accurately represent a wide variety of body shapes . It provides a function $M(\theta,\beta;\Phi):\mathbb{R}^{|\theta|\times|\beta|}\mapsto\mathbb{R}^{3N}$, mapping pose and shape parameters to $N=6,890$ vertices to form a mesh representation of human body. In the function $\theta\in\mathbb{R}^{24\times 3}$, $\beta\in\mathbb{R}^{10}$, and $\Phi$ are pose, shape, and model parameters, respectively. The pose parameters $\theta$ includes global rotation $\theta_{g}\in\mathbb{R}^{3}$ which is unrelated to the body pose and body pose parameters $\theta_{b}\in\mathbb{R}^{23\times 3}$, consequently $\theta=\{\theta_{g},\theta_{b}\}$. The shape parameters $\beta$ is coefficient of a low-dimensional shape space, learned from a training set of thousands of registered scans.
Typically, a pre-trained linear regressor $W$ is used to obtain $k$ keypoints of the whole body, i.e., $J_{3D}\in\mathbb{R}^{k\times 3}=WM(\theta,\beta;\Phi)$ [24]. As for 2D keypoints, a perspective camera model with fixed focal length and intrinsics $K$ is used to project 3D joint positions into 2D image plane. Each camera pose $\pi=\{R,t\}$ consists of a global orientation $R\in\mathbb{R}^{3}$ and a translation $t\in\mathbb{R}^{3}$. Given above parameters, 3D keypoints can be projected to 2D image as $J_{2D}=\Pi(K(RJ_{3D}+t))$, where $\Pi$ is a perspective projection with camera intrinsics $K$. Since the pose parameters $\theta$ already includes a global orientation $\theta_{g}$ which can represent the camera rotation, in practice we consider $R=\theta_{g}$, so when reconstructing human body from a single image, we estimate body pose parameters $\theta_{b}$ and camera pose $\pi=\{R,t\}$, and when reconstructing human body from $N$ images, we estimate body pose parameters $\theta_{b}$ and $N$ camera poses ${\pi_{i}=\{R_{i},t_{i}\}}_{i=1}^{N}$.
@@ 3.2 Unified Human Mesh Recovery (U-HMR)
Essentially, the multi-view feature fusion is a process of weighted sum of each view feature. In the transformer decoder structure, because of the cross-attention mechanism, the number of the weights is independent of the number of the input views. Due to this online variability, we can fuse features from arbitrary multi-view images. In addition, benefiting from the feature representation learning ability of the transformer decoder, the the updated SMPL query token can be more representative for human mesh recovery.
@@ 4.1 Experimental Setup
Human3.6M is a large 3D human pose benchmark, containing 11 different subjects and each subject with 15 different daily indoor actions. The subjects are captured by 4 synchronized 50 Hz digital cameras. The dataset provides 2D/3D keypoint annotations obtained by marker-based mocap system. The SMPL annotations are obtained by applying MoSh [36] to the sparse 3D MoCap marker following previous works [20, 25]. Similar to [20, 31, 47], we follow the protocol 1 to use subjects 1, 5, 6, 7, and 8 for training and subjects 9 and 11 are for testing.
MPI-INF-3DHP contains 8 actors performing 8 human activities. 2D and 3D keypoint annotations are also provided by marker-less mocap system. The dataset contains multi-view images of 14 cameras with regular lenses, and we choose views 0, 2, 7, 8 from them. Following [31, 47], the standard test dataset of MPI-INF-3DHP simply includes single-view images, we use subjects 1-7 for training and subject 8 for testing.
@@ 4.2 Ablation Studies
Table 2: Ablation study and comparison results of number of views on Human3.6M dataset.
@@
Figure 5: The performance trend w.r.t the number of views.
As we claim that our method can reconstruct human pose and shape for arbitrary multi-view images, to validate the flexibility, i.e., the ability to adapt to different number of views, we conduct ablation study on the effect of varying number of views during inference phase. Since there have been two methods, Yu et al. [61] and PaFF [19], state that they can be applied to arbitrary number of views, we compare with them on Human 3.6M dataset and the results are shown in Tab. 2. For fair comparison, our model is trained with a combination of varying number of multi-view image (1-view, 2-view, 3-view, and 4-view), which is same as  [61] does. While PaFF claims to have the ability to be applied to different number of views, the reported results are from the model retrained and tested using same number of views. Moreover, PaFF requires camera calibrations for feature fusion. So the results of PaFF can be expected to provide the performance upper bound of other two methods which are calibration-free and not retrained. Compared with Yu et al. [61], our method achieves better performance at any number of views.
We adopt varying number of views for training (1, 2, 4, 8) and inference (1-8) on MPI-INF-3DHP dataset. The performance trend w.r.t the number of views is in the Fig. 2. Note that in both training and inference phases the model precision grows with the increasing number of views and our model achieves reasonable performance when # training views $>$ # inference views or vice versa.
@@ 4.3 Comparison with State-of-the-art Methods
Table 3: Comparison results on Human3.6M dataset. We compare our method with single-view, calibration-requiring multi-view, and calibration-free multi-view methods. denotes the output of methods are non-parametric.
@@
4.4 Qualitative Comparison
Figure 6: Qualitative comparison results with HMR2.0 [11].
We show the results of comparison with HMR 2.0 [11] using images from indoor dataset, Human3.6M, and in-the-wild dataset, RICH, in the Fig. 6. It can be observed that our method provides more accurate body mesh (as shown in red boxes), and more importantly, the multi-view consistency.
5 Conclusion
Appendix 0.A Computational Overhead
@@ Appendix 0.B Sensitivity to Input
Table 7: Sensitivity to input. We perturb the input images of Human3.6M dataset and test model without retraining.
@@
Ours (Scale changing) 47.2 34.1 38.9 25.9
Ours (Center shifts) 52.6 35.6 41.7 26.7
@@
Appendix 0.C Images from Unseen Views
Figure 7: Visualization results of human mesh recovery from unseen camera views. We show the results from different number (1-4) of totally unseen views on MPI-INF-3DHPand TotalCapture datasets. For each sample, we show the input image, the reconstruction overlay and the top view.
Our method can recover human mesh from arbitrary multi-view images, including images from totally unseen camera views. For MPI-INF-3DHP and TotalCapture datasets, models are trained on 4 camera views by default as done in the main manuscript. As these two datasets provide more than 4 camera views, we test the model on totally unseen camera views of them, i.e., we randomly select an arbitrary number of unseen camera views and recover human mesh from them. The results are shown in Figure 7. We can observe that our method performs well even on totally unseen camera views.
Figure 8: Visualization results of human mesh recovery from arbitrary multi-view images. Results from 1-3 and 5-7 camera views on different datasets are shown. For each sample, we show the input image, the reconstruction overlay and the top view.
Figure 9: Visualization results of human mesh recovery from arbitrary multi-view images. Results from 4 and 8 camera views on different datasets are shown. For each sample, we show the input image, the reconstruction overlay and the top view.
Appendix 0.D Arbitrary Multi-view Images
In this section, we show visualization results of human mesh recovery from arbitrary multi-view images. As described in the main manuscript, the term “arbitrary” involves two characteristics: the arbitrary angle and arbitrary number of views. Thus, for each of the three datasets (Human3.6M, MPI-INF-3DHP and TotalCapture), we randomly select an arbitrary number of all available (seen and unseen) camera views and recovery human mesh from them. The results are show in Figure 8 and Figure 9. The results demonstrate the flexibility of our method.